# Learning How to Use Tools, Not Just When: Pattern-Aware Tool-Integrated Reasoning

### v2

@@ 1 Introduction
Tool-integrated reasoning (TIR) has become a powerful paradigm for enhancing large reasoning models (LRMs) wei2025autotir; chen2025r1; yu2025chain. By interacting with external verifiers such as code interpreters, TIR enables models to produce executable reasoning steps, making answers more accurate, faithful, and verifiable zheng2025deepresearcher. Recent progress has shown gains from better timing of tool invocation, multi-round tool calls, and tighter integration of external computation dong2025tool; wei2025autotir.
@@ 2.1 Reasoning with LRMs
LRMs have demonstrated better performance when explicitly generate intermediate reasoning stepsxu2025towards. Subsequent work has focused on improving the quality of these reasoning paths, either by introducing diversity through self-consistency wang2022self, refining intermediate steps via process supervision lightman2023let; jiang2025drp, or encouraging verification of intermediate results zelikman2022star; uesato2022solving.
@@ 2.2 Tool-Integrated Reasoning
Tool-integrated reasoning has emerged as a powerful paradigm for enhancing large reasoning models (LRMs) wei2025autotir; chen2025r1, enabling executable reasoning steps through external tools such as Python interpreters zheng2025deepresearcher. Early work like PoT chen2022program, PAL gao2023pal, and MathPrompter imani2023mathprompter showed that converting reasoning into code execution or lightweight snippets can substantially improve math performance, while implementations such as ToRA gou2023tora and Qwen2.5-Math-Instruct-TIR shao2024deepseekmath further demonstrated gains from deeper integration. However, prior studies primarily focus on when to call tools, overlooking how tools are applied once invoked. We address this gap by modeling pattern selection in TIR, introducing a pattern-aware framework that aligns tool-use strategies and yields significant improvements in reasoning accuracy.
@@ 3.3 Pattern Preference Alignment
We use DPO rafailov2023direct to empower our model with the ability to pick the correct pattern to solve math problems. For each problem $x$, we collect two candidate completions $y^{alg},y^{calc}$ and label a winner $y^{+}$ and loser $y^{-}$ using the teacher model. For a pair $(x,y^{+},y^{-})$, we then apply direction preference optimization with a frozen reference policy $\pi_{ref}$ (our SFT model):
@@ 4.1 Dataset
We use google Gemini-2.5-flash-lite as our teacher model, using our designed two pattern prompt to solve the problems in OpenR1-Math-220k openr1math220k dataset. From this corpus, we select the first 10,000 problems and divide them into training and validation sets using a 9:1 ratio. To assess complex mathematical reasoning and generalization, we evaluate on a broad set of out-of-domain benchmarks, including MATH500 math, AIME24 aime2024, and AMC23 amc2024.
@@ Appendix B Evaluate Pattern Chosen
To further examine the reliability of Gemini’s pattern selection, we asked GPT-5 to independently select patterns without prior exposure to Gemini’s choices. Across 100 test cases, GPT-5’s selections were consistent with Gemini’s in 98 instances, corresponding to a 98% agreement rate. This high level of concordance provides strong empirical support for the robustness of our approach. This follows LLM as a Judge li2025generation; gao2023human; jiang2024memorization method.

### v3

@@ 1 Introduction
Tool-integrated reasoning (TIR) has become a powerful paradigm for enhancing large reasoning models (LRMs) Wei et al. (2025); Chen et al. (2025); Yu et al. (2025). By interacting with external verifiers such as code interpreters, TIR enables models to produce executable reasoning steps, making answers more accurate, faithful, and verifiable Zheng et al. (2025). Recent progress has shown gains from better timing of tool invocation, multi-round tool calls, and tighter integration of external computation Dong et al. (2025); Wei et al. (2025).
@@ 2.1 Reasoning with LRMs
LRMs have demonstrated better performance when explicitly generate intermediate reasoning stepsXu et al. (2025). Subsequent work has focused on improving the quality of these reasoning paths, either by introducing diversity through self-consistency Wang et al. (2022), refining intermediate steps via process supervision Lightman et al. (2023); Jiang et al. (2025), or encouraging verification of intermediate results Zelikman et al. (2022); Uesato et al. (2022); Zhang et al. (2025).
@@ 2.2 Tool-Integrated Reasoning
Tool-integrated reasoning has emerged as a powerful paradigm for enhancing large reasoning models (LRMs) Wei et al. (2025); Chen et al. (2025), enabling executable reasoning steps through external tools such as Python interpreters Zheng et al. (2025); Jiang and Ferraro (2026). Early work like PoT Chen et al. (2022), PAL Gao et al. (2023a), and MathPrompter Imani et al. (2023) showed that converting reasoning into code execution or lightweight snippets can substantially improve math performance, while implementations such as ToRA Gou et al. (2023) and Qwen2.5-Math-Instruct-TIR Shao et al. (2024) further demonstrated gains from deeper integration. However, prior studies primarily focus on when to call tools, overlooking how tools are applied once invoked. We address this gap by modeling pattern selection in TIR, introducing a pattern-aware framework that aligns tool-use strategies and yields significant improvements in reasoning accuracy.
@@ 3.3 Pattern Preference Alignment
We use DPO Rafailov et al. (2023) to empower our model with the ability to pick the correct pattern to solve math problems. For each problem $x$, we collect two candidate completions $y^{alg},y^{calc}$ and label a winner $y^{+}$ and loser $y^{-}$ using the teacher model. For a pair $(x,y^{+},y^{-})$, we then apply direction preference optimization with a frozen reference policy $\pi_{ref}$ (our SFT model):
@@ 4.1 Dataset
We use google Gemini-2.5-flash-lite as our teacher model, using our designed two pattern prompt to solve the problems in OpenR1-Math-220k OpenR1 Team (2025) dataset. From this corpus, we select the first 10,000 problems and divide them into training and validation sets using a 9:1 ratio. To assess complex mathematical reasoning and generalization, we evaluate on a broad set of out-of-domain benchmarks, including MATH500 Hendrycks et al. (2021), AIME24 AI-MO Team (2024a), and AMC23 AI-MO Team (2024b).
@@ Appendix B Evaluate Pattern Chosen
To further examine the reliability of Gemini’s pattern selection, we asked GPT-5 to independently select patterns without prior exposure to Gemini’s choices. Across 100 test cases, GPT-5’s selections were consistent with Gemini’s in 98 instances, corresponding to a 98% agreement rate. This high level of concordance provides strong empirical support for the robustness of our approach. This follows LLM as a Judge Li et al. (2025); Gao et al. (2023b); Jiang and Ferraro (2024) method.